Money and Banking
Chapter 19
1. M = 2000, V = 5, Y = 20,000.
a. What does P equal?
b. What does k equal?
c. If M rises to 2500 and V and Y stay the same what does P equal? Show your work
d. Does an increase in V indicate that people want to hold more money or less?
e. Using the original values if V rises to 10 what happens to k?
f. Using the original values if V rises to 10 what does P equal? Show your work
g. If the growth of velocity = 0, the money supply growth = 5%, and output growth = 2%, what is the inflation rate according to quantity theory? Show your work.
1
Expert's answer
2020-04-03T10:26:14-0400

a)"P=\frac{M\times V}{Q}=\frac{2000\times5}{20 000}=0.5"


b)"k=\frac{M}{P\times Q}=\frac{2000}{0.2\times20 000}=0.5"


c)"P=\frac{M\times V}{Q}=\frac{2500\times5}{20 000}=0.625"


d) Yes, with an increase in the velocity of circulation, the economy develops and people become economically active, they want to have more money


e)"P=\frac{M\times V}{Q}=\frac{2000\times10}{20 000}=1"


"k=\frac{M}{P\times Q}=\frac{2000}{1\times20 000}=0.1"



f)"P=\frac{M\times V}{Q}=\frac{2000\times10}{20 000}=1"


g)"P=\frac{M\times V}{Q}=\frac{2100\times5}{20 400}=0.51"